I have been trying to solve these problems and I am unable to find the correct answers to them. If you can help me I would appriciate it.
Factoring
4c^2 -35c -9
56c+17c-3
s^2 +9sf+18f^2
v^2-7v+10
x^5y^5-x^4y^3+x^3y^2+x^2y^2
fv^2-27v-7
24c^2-18cd-42d^2
w^2-8w+16
12r^2-36r+27
25r^2-t^2
36a^2-100
(0.16y^2-0.0001)
9^12-16
b^3+125
16w^6-2f^6
1-a^8

Solving
0=w(w+2)
(2c-6)(6c-2)=0
v^2+2c-8=0
b^2+15b=0

The length of the top of a table is 6m greater than the width. The area is 72m. Find the demensions of the table.

This assignment is due today but I am willing to turn it in late. I really need the help and I would really appriciate it.

Answer by mananth(16946)

Factoring
4c^2 -35c -9
(c-9)(4c+1)
56c+17c-3
73c-3
s^2 +9sf+18f^2
(s+3)(s+6)

v^2-7v+10
(v-5)(v+2)
x^5y^5-x^4y^3+x^3y^2+x^2y^2
x^2y^2(x^3y^3-x^2y+3x+1)
fv^2-27v-7
???
24c^2-18cd-42d^2
6(4c^2-3cd-7d^2)
6(c(4c-7d)+d(4c-7d)
6(4c-7d)(c+d)
w^2-8w+16
(w-4)^2
12r^2-36r+27
3(4r^2-12r+9)
3(4r^2-6r-6r+9)
3(2r(2r-3)-3(2r-3))
3(2r-3)(2r-3)
3(2r-3)^2
25r^2-t^2
(5+t)(5-t)
36a^2-100
(6a+10)(6a-10)
(0.16y^2-0.0001)
(0.4y+0.01)(0.4y-0.01)
9^12-16
??
b^3+125
(b+5)(b^2-5b+25)
16w^6-2f^6
2(8w^6-f^6)
2((2w)^3-(f^2)^3)
2(2w-f^2)(4w^2+2wf^2+f^4)
1-a^8
(1+a^4)(1-a^4)
(1+a^4)(1-a^2)(1+a^2)
(1+a^4)(1+a^2)(1+a)(1-a)

Solving
0=w(w+2)
w= 0 OR w=-2
(2c-6)(6c-2)=0
c=3 OR c=1/3

v^2+2c-8=0
??
b^2+15b=0
b(b+15)=0
b=0 OR b=-15

The length of the top of a table is 6m greater than the width. The area is 72m. Find the demensions of the table.

This assignment is due today but I am willing to turn it in late. I really need the help and I would really appriciate it.